### v1

# Chatbots and Zero Sales Resistance

Abstract
It is argued that the pursuit of an ever increasing number of weights in large-scale machine learning applications, besides being energetically unsustainable, is also conducive to manipulative strategies whereby Science is easily served as a strawman for economic and financial power. If machine learning is meant to serve science ahead of vested business interests, a paradigm shift is needed: from more weights and little insight to more insight and less weights.
1 Introduction
Not a day goes by without we hear of the latest AI breakthroughs, such as chatbots that write up texts or generate images increasingly harder to tell apart from their human-made counterparts. These headlines come with a heavy load of hype, but even with hype factored out, a highly seductive promise stands tall, the promise to capture levels of complexity largely out of grasp for our best theories, models and simulations. Briefly, AI would supplant the time-honored Scientific Method, as we know it since Galileo’s time [1, 2].
While heavily pumped up, this promise is not empty, addressing as it does, among others, one of the most vexing Achille’s heels of the scientific method, the infamous Curse of Dimensionality (CoD) [3]. Indeed, CoD compounds with a profound hallmark of Complexity, namely the fact that complex systems are sneaky: they inhabit ultra-dimensional spaces but don’t fill them up [4, 5, 6]. To the contrary, ”interesting things” take place in ultrathin and often highly scattered portions of the huge state space available to them. Nature likes to play hide and seek and big time so. An illuminating example can be found in the book of Frenkel and Smit [7], where we learn that the chance of making a sensible Monte Carlo move in the state space of hundred hard-spheres (please note, hundred, not Avogadro’s) is about $10^{-260}$! The golden nuggets are well hidden indeed.
Computational science has devised a number of clever techniques to visit the regions hosting the preciously rare golden nuggets without waiting many ages of the Universe [8]. Yet, the CoD still remains a very tough cookie for the scientific method to the present day.
Artificial Intelligence, and more specifically Machine Learning (ML), promise a new and unprecedently powerful angle of attack to Complexity. And again, the promise is largely overblown but not empty, as witnessed by a number of success stories: chess and GO winnings, self-driving cars, DeepFold mapping of protein structure, stand out as some the most spectular(ized) cases in point [9].
2 The basic ML procedure
It is worth discussing where this ”magic” comes from in some little more detail.
The basic idea of ML is to represent a given $D$-dimensional output $y$ (target) through the recursive application of a simple non linear map [10]. For a neural network (NN) consisting of an input layer $x$, $L$ hidden layers $z_{1}\dots z_{L}$, each containing $N$ neurons, and an output layer $y$, the update chain $x\to z_{1}\dots\to z_{L}\to y$ reads symbolically as follows:
$z_{1}=f(W_{1}x-b_{1}),\;\dots z_{L}=f(W_{L}z_{L-1}-b_{L}),\;y=f(W_{L+1}z_{l}-b_{L+1})$ (1)
where $W_{l}$ are $N\times N$ matrices of weights, $b_{l}$ are N-dimensional arrays of biases and $f$ is a nonlinear activation function, to be chosen out of a large palette. The output $y$ is then compared with a given training target $y_{T}$ (Truth) and the weights are recursively updated in such a way as to minimize the discrepancy between $y$ and $y_{T}$ (Loss function), up to the desired tolerance. In equations
$\mathcal{L}[W]=dis(y[W],y_{T})\leq\epsilon$ (2)
where the loss function $\mathcal{L}$, namely the distance between the model output and the target in some suitable metric, is an explicit function of the weights configuration $W$. The central engine of the ML procedure is an update schedule of the weights whereby the loss function is taken below the desired tolerance $\epsilon$. This is typically obtained through some form of steepest descent (SD) search:
$W(t+1)=W(t)-\gamma\frac{\partial\mathcal{L}}{\partial W}$ (3)
where $t$ denotes the iteration step and $\gamma$ is a suitable relaxation rate.
The bold idea is that with big enough data for training, the ML sequence (1-2-3) can reach any target, whence the alleged demise of the scientific method [1, 2].
Where does such magic come from?
The key point is that for a DNN (deep neural net) of depth $L$ (number of layers) and width $N$ (numbers of neurons per layer), there are $N_{P}=N^{L}$ possible paths connecting any single item $x_{i},\;i=1,N$ in the input layer to any another single item $y_{j},\;j=1,N$ in the output layer. Hence a DNN with $N=10^{3}$ neurons and $L=10^{2}$ layers features $N_{W}=N^{2}L=10^{8}$ weights and $N_{P}=10^{30}$ paths. Such gargantuan network of paths represents the state-space of the DNN learning process, which can proceed through several concurrent paths at a time. If you think that this is sci-fi, please think again, as current leading edge ML applications, such as DeepFold or Large Language Models motoring the most powerful ”ask-me-anything” chatbots are using up to 100 billions weights, basically the number of neurons in our brain. Except that our brain works at 20 Watt while the largest ML models are now sucking up at least ten million times more, a point to which we shall return shortly.
These numbers unveil the magic behind ML: DNN duel the CoD face up, by unleashing an exponential number of paths, and adjusting them in such a way as to sensibly populate the sneaky regions where the golden nuggets are to be found.
This strategy is an opinion splitter: AI pragmatists are enthusiastic at the conceptual simplicity of this black-box and, speaking of weights, to them ”too much is not enough”. Scientists fond of Insight long ahead of Control, are horrified at the diverging number of parameters, their ”prejudice” being that parameters are fudge factors concealing lack of understanding, so that their motto is rather ”the least the best”.
3 Criticism to ML
There are indeed several solid reasons to second-guess the blue-sky scenarios portrayed by the most enthusiastic (and aggressive ) AI aficionados. Here we briefly summarize the main items, some of which have been pointed out before [11, 12]:
1) Convergence: the convergence of the ML procedure depends on the landscape of the Loss function in weight space, meaning that the SD procedure may get trapped in false local minima. This is particularly true in the case of hyperdimensional weight spaces.
2) Accuracy: in order to achieve the required level of accuracy enormous data sets might be required. This is because uncertainty in complex systems is far more resilient than the standard gaussian rule $1/\sqrt{N}$, $N$ being the size of the data set. The reason is physical: the gaussian rule applies to uncorrelated systems, whereas complex systems are typically highly correlated, in both space and time.
3) Explainability: the optimal weights, $W_{opt}$, associated to a given solution, do not lend themselves to any direct physical interpretation. In other words, they provide no Insight into the phenomenon under study.
4) Causality: the ML procedure is built upon the idea of catching Correlations between the model solution $y=f(x,W)$ and the target $y_{T}$. As is well known, this says nothing about the causal relations that underlie such correlation. Once again, Control without Insight. Moreover it is well known that the ratio between false and true correlations is an exponential function of the size of the dataset [13].
5) Uncertainty Quantification: once an optimal weight configuration $W_{opt}$ is found, it is extremely hard to estimate the robustness of the solution toward changes of the weight parameters. The main hurdle is again the hyperdimensionality of weight space.
6) Smoothness: the activation functions commonly used in ML are differentiable. While expedient to the simplicity of the SD procedure, this hampers convergence in case the output is a rough (non-differentiable) function of the input, as it is often the case in complex systems.
7) Sustainability: Training does not come for free: it is estimated that new generations of chatbots with trillions of weights may reach up to Gigawatts of power demand.
The above list of items is open for a healthy and hopefully hype-free scientific debate. Yet, it appears hard to argue against the idea that a blind escalation of the number of weights is hardly the best way to go, on both scientific and societal accounts.
In the following we focus on latter, with special emphasis on issue 7) above, which is clearly tightly intertwined with all other items in the list.
4 Sustainability
Besides the exponential number of parameters, in order to converge, the ML procedure needs correspondingly huge training datasets. As mentioned above, training does not come for free: it is estimated that new generation chatbots will come near to the Gigawatt power demand, in excess of most existing power plants. The comparison with the 20 Watts of our brain is embarrassing. True, once trained, their routine operation is much less expensive. The question still remains, as to whether, chatbots can still function accurately and reliably on unseen data.
Given the energy-devouring nature of the large-scale ML procedure, the obvious question is: is it really worth sacrificing a substantial share of the total energy budget worldwide to the totem of chatbots?
This may make Jensen Huang, the founder of Nvidia, the richest man on Earth, but still it is not the way to go for the rest of us, apart the super-elite who may be able to afford extra-terrestrial life for the decades to come. The point is that, even when it works, ML is hardly Explainable, it offers little clue on the physical meaning of the parameters: Control but little Insight, if any at all. And with Insight out of the game, there’s no guarantee that what works for seen data will keep working again for the unseen ones (extrapolation).
Of course, one can close eyes and keep going on steroids with weights, but, given the power bill discussed above, this sounds reckless at best. There must be better ways. An interesting clue in this direction is that the overwhelming majority of the weights, the experts tell me, are close to zero, meaning that they do not contribute significantly to the success of the ordeal, hence offering little return on computing investment. This is scientifically very interesting and it begs for understanding, not just because this is what science is all about, but also because Understanding here means saving oceans of Gigawatts.
Recently Elon Musk advocated the need of putting AI on a rigorous scientific basis, in his own words ”Join xAI (Explainable AI) if you believe in our mission of understanding the universe, which requires maximally rigorous pursuit of the truth, without regard to popularity or political correctness.” Yann LeCun, one of the most respected computer scientists worldwide and Turing awardee, promptly countered that ”Musk wants a maximally rigorous pursuit of the truth but spews crazy-ass conspiracy theories on his own social platform.” And, upon being questioned by Musk about his recent science, LeCun goes on by quoting his some 80 papers, as opposed to Musk’s zero entries in this ballgame.
Now, it is ironic enough to hear one of the most muscular and hungriest AI energy consumers on the planet to advocate the rigorous pursuit of scientific truth. And despite his towering status in computer science, it is only slightly less ironic to see LeCun taking on the role of the guardian of science, given that LeCun, incidentally also Chief AI Scientist at Zuckerberg’s Facebook, is a champion of that kind of computer science where the dismissal of Insight in favor of Control is largely tolerated, see [14, 15].
But let’s give Musk the benefit of doubt and assume he’s genuinely interested in understanding how the Universe works. The tip is fairly easy: stop leveraging the muscular power of ML with legions of GPU’s and surprise us with more understanding and less weights. The name of the game being causal AI or explainable AI, acknowledging that, despite the wildest claims of AI aficionados, whenever Correlation can replace Causality, what we are talking about is not Science but Control.
This is the biggest lie which has been served to us by AAI, where AAI stands for Aggressive AI, in order to distinguish it from the many important contributions of AI to technology and society. But the king is now naked, for the pursuit of Control regardless of Insight comes with a energy price tag that planet Earth just cannot sustain. Incidentally, scientists (most physicists and mathematicians) much less visible than Musk and LeCun, are already doing this, out of the limelight [16, 17].
5 Zero sales resistance
Musk’s glorious statement about xAI calls for another comment, mostly related to a strategy that, back in 1944 in his prophetic book ”The abolition of man”, CS Lewis called ”Zero Sales Resistance”, ZSR for short [18]. Pursuing financial/economic interests under the glorified veil of world-saving intentions is a well-known strategy since long: what is new, though, is the unprecedented power of modern AAI. Here comes the point. The speed of modern technology has basically collapsed spacetime: a single message can reach any (connected) individual on this planet in virtually no time. The result is that a well-crafted message, suitably conveyed by your most seductive influencer of choice (maybe a chatbot?), can win billions of brains in a single swoop, a process called brain condensation [19]. The associated profits go with some square root of this condensation ratio, as reflected by the four-five orders of magnitude gap between the salary of top executives versus their least paid employees [20]. It is perhaps no coincidence to hear Zuckerberg proclaiming that ”connectivity is a human right”. The next interesting step is to realize that steering the ”sentiments” of human beings amounts to controlling a comparatively small number of high-level ”psychological variables”, a piece of cake for the most powerful ML applications, with little or no need of Insight [19].
Trying to cure Alzheimer with ML shows a very different movie, one where CoD hits hard, Correlation cannot replace Causation, Control cannot replace Insight, a ballroom called Science [21]. Yet, the tools are the same, it is basically the very same machinery described by the equation (1) above! That’s why Science is so easily served as a strawman for Control.
This is where AAI becomes as dangerous as never before in the history of science: it just trickles into our habits, step by step, no trauma, no shock, a silent but relentless conquer of our brain towards the ZSR goal. And please, make no mistake, the main item on sale with ZSR is not the commercial product but our brain instead.
The motto Why learn if you can look it up? speaks loud for the above.
So, while science eventually gets important contributions from ML, if only hardly in proportion to bombastic claims, sky-rocketing sales behind the scenes is a sure thing. CS Lewis says it best: ”When men over-emphasize the control over nature as scientific progress, what they really mean is control of a few human beings over many other human beings”.
6 A call to algorethics
What to do then?
In the recent years there have been increasing calls to ”Algorethics”, the idea being to inject ”ethical constraints” into the ML algorithms [22]. This is certainly a commendable goal and, as nicely discussed in Kearn and Roth’s book [23], quite likely a technically doable one as well.
Question is: will it really fix the ZSR issue?
I sincerely doubt it, no matter how good the law, history shows that the smart villain always manages to find the next loophole. More effective, I think, is to pursue the inherent spiritual drive of Science (and capital R Religion, of course), namely the pursuit of Insight for the pleasure of finding things out, period. The rest will follow, it always did.
7 Summary
Machine learning has potential to provide a substantial new entry to the world of scientific investigation, as it can occasionally capture levels of complexity beyond reach for our best theoretical and computational models. However the blind pursuit of ever-increasing number of weights with little or no attention to Insight versus Control, is overly dangerous in several respects. Besides being energetically unsustainable, it paves the way to zero-sales-resistance strategies whereby Science is served as a perfect strawman for Control and financial power.
8 Acknowledgements
The author is grateful to SISSA for financial support under the ”Collaborations of Excellence” initiative as well as to the Simons Foundation for supporting several enriching visits of his. He also wishes to acknowledge many enlightening discussions over the years with PV Coveney, A. Laio and D. Spergel. I am also grateful to M. Durve for critical reading of the manuscript.

### v2

@@
 3 Criticism to ML
-There are indeed several solid reasons to second-guess the blue-sky scenarios portrayed by the most enthusiastic (and aggressive ) AI aficionados.
+There are indeed several solid reasons to second-guess the blue-sky scenarios portrayed by the most enthusisatic (and aggressive ) AI aficionados.
 Here we briefly summarize the main items, some of which have been pointed out before [11, 12]:
@@
 As mentioned above, training does not come for free: it is estimated that new generation chatbots will come near to the Gigawatt power demand, in excess of most existing power plants.
-The comparison with the 20 Watts of our brain is embarrassing.
+The comparison with the 20 Watts of our brain is embarassing.
 True, once trained, their routine operation is much less expensive.
@@
 But the king is now naked, for the pursuit of Control regardless of Insight comes with a energy price tag that planet Earth just cannot sustain.
-Incidentally, scientists (most physicists and mathematicians) much less visible than Musk and LeCun, are already doing this, out of the limelight [16, 17].
+Incidentally, scientists (most physicists and mathematicians) much less visible than Musk and LeCun, are already doing this, out of the limelight [16, 17, 18].
 5 Zero sales resistance
@@
 However the blind pursuit of ever-increasing number of weights with little or no attention to Insight versus Control, is overly dangerous in several respects.
-Besides being energetically unsustainable, it paves the way to zero-sales-resistance strategies whereby Science is served as a perfect strawman for Control and financial power.
+Besides being energetically unsustainable, it paves the way to zero-sales-resistence strategies whereby Science is served as a perfect strawman for Control and financial power.
 8 Acknowledgements